th equation of tangent to the circle x sqr+ysqr=25 passing through (-2,11)

Grade:

1 Answers

Arpit Kumar
10 years ago
The equation of tangent to a circle having equation of the form x^2+y^2=a^2 and passing through the point x1,y1 is given by xx1+yy1=a^2 so the equation of the tangent will be 11y-2x=25 Hope this helps